OpenStudy (anonymous):

Central Middle School has 2,470 students and the ratio of students to teachers is 26:1 How many new teachers are needed to reduce this ratio to 19:1?

OpenStudy (anonymous):

@HomeschoolGrad can you help me please

OpenStudy (anonymous):

Here I got dis

OpenStudy (anonymous):

set up your proportion first

OpenStudy (anonymous):

2470/x=26/1 cross multiply 2470=26x 2470/26=x 95=x

OpenStudy (anonymous):

now to change it to 19:1

OpenStudy (anonymous):

2470/x=19:1 cross multiply again 2470=19x 130=x

OpenStudy (anonymous):

You have 95 teachers and need 130 130-95=35 you need 35 new teachers

OpenStudy (anonymous):

so i just divide 2470 to 26 and get 95 and for the second one i divide 2470 to 19?

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

You're welcome:) But do you understand why you divide that? becaus of proportions

OpenStudy (anonymous):

yes I had the right answer but i was just making sure i was doing it correctly

OpenStudy (anonymous):

ok then yes